OpenStudy (samara89):

Adanna had a glass cylinder with a radius of 6 cm and a height of 25 cm. She placed 3 balls, each with a diameter of 6 cm, into the cylinder. She filled the rest of the cylinder with colored water. How many cubic centimeters of colored water did she use? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (samara89):

@greatlife44

OpenStudy (samara89):

I got 2486.88cm^3 is that correct

OpenStudy (samara89):

@welshfella

OpenStudy (greatlife44):

\[\pi*r^2*h_{cylinder~volume} = (6cm)^{2}*(25cm)*\pi = 900\pi_{volume~cylinder}\] \[\frac{ 4 }{ 3 }(3cm)^{3}*\pi*3 = 339cm^{3}_{~balls}\] \[(900\pi-339)cm^{3} = 2487cm^{3}_{water}\]

OpenStudy (greatlife44):

nice work @Samara89

OpenStudy (samara89):

so the answer is 2487

OpenStudy (welshfella):

Yes Good work

OpenStudy (greatlife44):

you know what you're doing